6. If a company sets aside $1,000,000 now into a contingency fund, how much will the company have in 2 years, if it does not use any of the money and the account grows at a rate of 10% per year Compounded ?
Answer #1

Answer : Here

Principal (P) = $1,000,000

Rate of interest (r) = 10% = 0.1

Time (t) = 2 years

n = Per year interest rate compounding time = 1

Now,

Amount = P (1 + r/n)^nt

=> Amount = 1,000,000 (1 + 0.1/1)^(1*2)

=> Amount = 1,000,000 (1.1)^2

=> Amount = 1,000,000 * 1.21

=> Amount = $1,210,000

Therefore, in 2 years the company will have $1,210,000.